### v1

@@ 2.2 Spectroscopic data
We use the spectroscopic observations of HCO+(43) line at 356.734 GHz obtained from the Heterodyne Array Receiver Program (HARP; Buckle et al., 2009) mounted on JCMT. The observations were done during June and July 2011 (Proposal ID: M11AU13), covering the Oph-A region with a beam size (or FWHM) of 14. We use the reduced spectral map of HCO+(43) line directly downloaded from the JCMT science archive222https://www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/en/jcmt/. The data cube has an original spectral resolution of 0.025 km s-1 and is smoothed to have a final spectral resolution of 0.5 km s-1. For comparison to the HCO+(43) data, we also use observations of N2H+(32) line at 279.515 GHz obtained from the APEX-2 instrument on the 12-m Atacama Pathfinder EXperiment telescope (APEX; Güsten et al., 2006). The data cube has an original spectral resolution of 0.082 km s-1. Here, we use the back-end data cube with a spectral resolution of 0.25 km s-1, covering the Oph-A region with the beam size of 22 (for details, see Liseau et al., 2015).
@@ 2.3 Dust temperature and column density maps
We adopt the maps of dust temperature () and gas column density () from Santos et al. (2019). These maps were obtained from the fitting of a modified black body to the spectral energy distribution (SED) using the continuum flux maps from Herschel/PACS (Poglitsch et al., 2010) at 70, 100, and 160 m and have the FWHM of 11.4 (for details, see Section 3.2 in Santos et al., 2019). Figure 1 shows the spatial distribution toward Oph-A, overlaid with contours of column density . In this study, we will use the map of as the map of the gas kinetic temperature, assuming that the gas is in the local thermodynamic equilibrium (LTE) and well-coupled with the dust.
@@ 3.1 Magnetic field morphology
We construct the B-field morphology in Oph-A by rotating polarization angles by 90°. Figure 2 shows the maps of B-field orientation (half-vectors) inferred from maps of the polarization angle at 89 and 154 m . The background color maps show the Stokes- intensity at the respective wavelength. Here, the position angle of B-fields () is measured east of north, and in the range between 0 and 180. The polarimetric data at both wavelengths can well reveal B-fields toward the dense region in Oph-A where Stokes-¿0.12 Jy/arcsec2 (the area inside the lowest contour in each map). The data at 89 m can further resolve the B-fields morphology extended toward the warmer dust region located at the east side of the Oph-S1 star.
The B-field vectors in Oph-A are generally well-ordered and largely follow the material structure in low- to intermediate-densities regimes (see Figure 2). B-field orientations seem to be bent toward the higher-density region in the center of the map and are mostly perpendicular to the ridge of the cloud. On the eastern side of the map, the orientation of B-field vectors tends to vary in the counter-clockwise direction from North to South, i.e., the position angle of B-fields increases. Particularly, in the North-East part of the map, B-field vectors run along the North-East to South-West direction with a median value of 56° and 49° at 89 and 154 m, respectively. This feature is also seen in a small area in the North-West part of the map. Toward the East side, B-field vectors orient nearly to the East-West direction, with median values of 83° at 89 m and 75° at 154 m. On the other hand, B-field vectors in the South-East region tend to orient consistently to the South-East to North-West directions, with median values of 92° and 85° at 89 and 154 m, respectively. Interestingly, the data at 154 m shows B-field vectors running nearly vertically, i.e., along the North-South direction toward a small area at the North part of the map, which is not seen in the data at 89 m (see Figure 2 and also in Figure 15).
In addition, there are two outstanding components of B-fields toward the regions characterized by the highest H2 column densities: one component associated with the peak of continuum emission at 154 m runs along North-East to South-West direction with the median B-field orientation of 70; the other is located at the North of the first component and associated with the peak of continuum emission at 89 m, which exhibits the B-field vectors along the East-West direction with a median position angle of 88° and 81° at 89 and 154 m, respectively (see Figure 15). We will further discuss the B-field morphology together with B-field strengths in Section 4.1.
@@ 3.2 Measurement of magnetic field strength using the DCF method
We first use the classical DCF method (Davis, 1951; Chandrasekhar & Fermi, 1953) (hereafter we refer to as the DCF method) to calculate the strengths of the B-field projected in the plane-of-the-sky () in the entire region of the Oph-A cloud. We will compare our obtained results with those obtained using more advanced techniques in Section 3.3.
@@
where is the correction factor, is the gas mass density in units of g cm-3, =2.8 is the mean molecular weight per hydrogen molecule (Kauffmann et al., 2008), is the mass of the hydrogen atom, is the gas volume density in units of cm-3), is the one-dimensional (1D) non-thermal velocity dispersion along the line-of-sight of the ideal gas species that traces the dense region as similar as by dust polarized emission in units of km s-1, and is the polarization angle dispersion in degrees. The correction factor is uncertain, and numerical simulations by Ostriker et al. (2001) show when the polarization angle dispersion does not exceed 25. However, varying values of will result in different strengths (see e.g., Skalidis et al., 2021; Liu et al., 2021, for detail of discussions).
Here, we use the formula as appeared in Crutcher (2004):
@@ 3.2.1 Polarization angle dispersion maps
It is essential to see how the B-field orientation changes locally within the molecular cloud. To do that, we calculate the dispersion in the polarization angle distribution () across the Oph-A region. We use the method presented in Hwang et al. (2021) for which is calculated pixel-by-pixel throughout the map of Oph-A. We summarize the procedure of calculating toward the pixel in the map as follows:
@@
We work with a box covering the pixel with a size of 99 pixels, corresponding to 2 beam sizes of the SOFIA/HAWC observations (i.e., 15.6 and 36.4 in bands C and D, respectively).
@@
We calculate the mean value of polarization angles within the small box, , which implies the local mean polarization angle toward the pixel .
@@
For each pixel within the small box, we calculate the difference between the observed polarization angle and the mean value. We compute the root mean square (RMS) of the distribution of angle differences in the box. This RMS value indicates the polarization angle dispersion calculated toward the pixel .
@@
By repeating the process for all pixels in the entire region, we obtain the map of in Oph-A. We note that if there are more than 50 of pixels undefined (i.e., having NaN value of ) in the computing box, the result of polarization angle dispersion calculation toward pixel is ignored. We apply this criterion to keep the high quality of the estimate. It is also necessary to de-bias the polarization angle dispersion due to the uncertainty of the observed polarized angles.
Figure 3 shows the maps of polarization angle dispersion toward Oph-A using HAWC data in two bands C (left panel) and D (right panel), respectively. The polarization angle dispersion calculated using data in bands C and D are in ranges 2.3–32.2 and 1.3–31.5, which correspond to the median values of 5.7° and 5.8°, respectively. The differences of in two wavelengths imply that the data at the shorter wavelength tends to trace warm dust surrounding the high-mass star S1 and toward the outer layers; whereas the data at the longer wavelength probes the colder dust in the denser layers of the cloud. We note, however, that in maps of in both bands, there are a few pixels where (27 pixels in band C and 45 pixels in band D). Thus, we do not use the value in these pixels to estimate in Section 3.2.4.
@@ 3.2.2 Velocity dispersion map
We use the HCO+(43) molecular data observed by JCMT/HARP to estimate the total velocity dispersion, . This molecular line has a critical density of 106 cm-3 (Greve et al., 2009) and thus is considered a good tracer to probe the dense region in the Oph-A cloud. The HCO+(43) emission is well-represented by a single Gaussian profile. Thus, in each pixel, we fit the line with a single Gaussian. The central velocity () across Oph-A covers the range from 2.25 to 3.76 km s-1 with a median value of 3.12 km s-1, which is in agreement with the systemic velocity of Oph-A of 3.44 km s-1 (André et al., 2007). We use the line width of the Gaussian as the measure of total velocity dispersion toward each pixel.
To construct the map of the 1D non-thermal (turbulent) velocity dispersion (), we subtract the thermal component () from the map of the total velocity dispersion, given by:
where =29 amu is the mass of the HCO+ molecule, is the Boltzmann constant, and is the kinetic temperature of the gas. Here, we assume that the gas is in LTE condition and well-coupled with the dust in the cloud. Thus, the gas kinetic temperature is equal to the excitation temperature in all levels () and equal to the dust temperature: . We convolve the map of to the lower resolution of the molecular line map (14) to obtain the same pixel size on both maps. We convert the velocity dispersion of the non-thermal component into the FWHM velocity dispersion, given as . The uncertainty of non-thermal FWHMs is propagated from the uncertainty of the line width obtained from the Gaussian fit.
Figure 4 shows the map of the 1D non-thermal FWHM of the HCO+ molecular line toward Oph-A. The values are in the range from 0.62 to 2.98 km s-1 with a median value of 1.19 km s-1. We find a good agreement between our result obtained with HCO+(43) and those from other line tracers, i.e., 0.95 km s-1 using N2H+(32) from APEX and 1.5 km s-1 using C18O(32) from JCMT/HARP (see Appendix C for details). Thus, it is plausible to use results from the HCO+(43) line in the following analysis.
@@ 3.2.3 Gas volume density map
We estimate the gas volume, , across the dense cloud from . Here, we assume that the cloud has a spherical Plummer-like shape (Liseau et al., 2015; Santos et al., 2019), which is frequently used to describe the globular dense cores (Plummer, 1911). We consider a sphere centered at the peak position of , located at (RA, Dec)= (, ), which is also associated with the position of the starless core SM1 (the separation is much less than 1″). We take the radius of the cloud (0.074 pc, Santos et al., 2019). To describe the effect of stellar feedback from the high-mass Oph-S1 star, we further add a spherical shell with a radius of surrounding this star. We also assume that the positions of peak column density and the S1 star are located in the same plane perpendicular to the line of sight. The top panels in Figure 5 show the 3D schematic view of the cloud (left) and a cross-section on the -plane at a given position (right). The - and -axes perpendicular to each other are in the plane of the sky and the axis is along the line of sight. The center of the sphere is located at . The depth of the cloud at each position () along axis is calculated by:
where and are the radius of horizontal circles perpendicular to the plane of sky at position (see top right panel in Figure 5). We obtain the map of , shown in the bottom left panel in Figure 5. The values of are in the range of . We then infer from using the formula:
The bottom right panel in Figure 5 shows the spatial extent of across Oph-A, overlaid by the contours representing the continuum emission in band D (Stokes-). The map has an angular resolution of 11.4″and a pixel size of 3.2, similar to that of the map. The gas density spans from cm-3 and is highest at the center of the cloud. The pattern of distribution well follows the intensity map in band D, revealing the ridge shape of the dense cloud. Some pixels located at the circular boundary in the map show anomalously high densities; this is due to the very low depths caused by the spherical model, and thus we exclude these pixels from our strength measurement.
@@ 3.2.4 Magnetic field strength maps
We use maps of polarization angle dispersion (), non-thermal FWHM velocity dispersion (), and gas volume density () to measure in the entire region of Oph-A. All input maps are smoothed to the angular resolution of 14, corresponding to that of the map. We then reconstruct the maps of at 89 and 154 m toward Oph-A using Eq. 2, which are shown in Figure 6.
The values vary from mG and from mG for data at 89 and 154 m, respectively. The highest strength is found at the center region of the cloud, associated with the emission peak at 154 m and also the position of the SM1 core. The median values are and mG333The uncertainties of these median values are solely of the estimated distribution across the cloud. , as measured using 89 and 154 m data, respectively. The uncertainty of is propagated based on the uncertainty of the velocity dispersion of the gas and polarization angle dispersion. Since map is inferred from resulted in SED fitting models (see Santos et al., 2019). In addition, the estimated depth of the cloud is based on the 3D geometrical shape of the cloud and also is uncertain. Thus, we do not take into account the uncertainty of the gas volume density in the total uncertainty of the B-field strengths. For this reason, the uncertainty of might be underestimated. Due to the lack of measurements of B-field strengths along the line-of-sight () in the Oph-A cloud, we adopt the component as the total B-field strengths : . For the following analysis, we will employ only the map inferred from the polarimetric data at 154 m since it covers the larger area of the higher-density region in Oph-A than that from the data at 89 m.
@@ 3.3 Measurement of magnetic field strength using other methods
Several new techniques have been proposed to improve the DCF method for the last decades. For instance, Hildebrand et al. (2009) introduced the angle dispersion function (ADF) method where the dispersion in polarization angles in Eq. 2 is replaced by the square root of the structure function of polarization angles. This provides an improved measurement of B-field strength but also gives us the ratio of the turbulent to large-scale B-field strengths. Alternatively, Skalidis & Tassis (2021) introduced a modification of the DCF method, hereafter ST, assuming that the mean (ordered) and fluctuation components of the total B-field energy, and , respectively, are not independent. More recently, the differential measure approach (DMA; Lazarian et al., 2022) is considered promising to enhance the accuracy of B-field strength measurements at scales smaller than the turbulence injection scale and takes into account the anisotropic properties of MHD turbulence (given by factor ). For the sake of cross-checking the B-field strength measurement in Oph-A using the DCF method, we re-calculate the mean B-field strengths using ADF, ST, and DMA methods at the two subregions (a) and (e) associated with the densest parts of the cloud (these subregions are named as in Kwon et al. (2018), see also in Figure 15).

### v2

@@ 2.2 Spectroscopic data
We used the spectroscopic observations of HCO+(43) line at 356.734 GHz obtained from the Heterodyne Array Receiver Program (HARP; Buckle et al., 2009) mounted on JCMT. The observations were carried out during June and July 2011 (Proposal ID: M11AU13), covering the Oph-A region with a beam size (or FWHM) of 14. We used the reduced spectral map of HCO+(43) line directly downloaded from the JCMT science archive222https://www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/en/jcmt/. The data cube had an original spectral resolution of 0.025 km s-1 and was smoothed to have a final spectral resolution of 0.5 km s-1. For comparison to the HCO+(43) data, we also used observations of N2H+(32) line at 279.515 GHz obtained from the APEX-2 instrument on the 12-m Atacama Pathfinder EXperiment telescope (APEX; Güsten et al., 2006). The data cube has an original spectral resolution of 0.082 km s-1. Here, we used the back-end data cube with a spectral resolution of 0.25 km s-1, covering the Oph-A region with the beam size of 22 (for details, see Liseau et al., 2015).
@@ 2.3 Dust temperature and column density maps
We adopted the maps of dust temperature () and gas column density () from Santos et al. (2019). These maps were obtained from the fitting of a modified black body to the spectral energy distribution (SED) using the continuum flux maps from Herschel/PACS (Poglitsch et al., 2010) at 70, 100, and 160 m, with an FWHM of 11.4 (for details, see Sect. 3.2 in Santos et al., 2019). Figure 1 shows the spatial distribution toward Oph-A, overlaid with contours of column density, . In this study, we use the map of as the map of the gas kinetic temperature, assuming that the gas is in the local thermodynamic equilibrium (LTE) and well-coupled with the dust.
@@ 3.1 Magnetic field morphology
We construct the B-field morphology in Oph-A by rotating polarization angles by 90°. Figure 2 shows the maps of B-field orientation (half-vectors) inferred from maps of the polarization angle at 89 and 154 m . The background color maps show the Stokes- intensity at the respective wavelength. Here, the position angle of B-fields () is measured east of north, and in the range between 0 and 180. The polarimetric data at both wavelengths can well reveal B-fields toward the dense region in Oph-A, where Stokes-¿0.12 Jy/arcsec2 (the area inside the lowest contour in each map). The data at 89 m can further resolve the B-fields morphology extended toward the warmer dust region located at the east side of the Oph-S1 star.
The B-field vectors in Oph-A are generally well-ordered and largely follow the material structure in low- to intermediate-densities regimes (see Fig. 2). B-field orientations seem to be bent toward the higher density region in the center of the map and are mostly perpendicular to the ridge of the cloud. On the eastern side of the map, the orientation of B-field vectors tends to vary in the counter-clockwise direction from north to south; thus, the position angle of B-fields increases. Particularly, in the northeast part of the map, B-field vectors run along the northeast to southwest direction with a median value of 56° and 49° at 89 and 154 m, respectively. This feature is also seen in a small area in the northwest part of the map. Toward the east side, B-field vectors orient nearly to the east-west direction, with median values of 83° at 89 m and 75° at 154 m. On the other hand, B-field vectors in the southeast region tend to orient consistently to the southeast to northwest directions, with median values of 92° and 85° at 89 and 154 m, respectively. Interestingly, the data at 154 m shows B-field vectors running nearly vertically, that is, along the north-south direction toward a small area at the north part of the map, which is not seen in the data at 89 m (see Fig. 2 and also Fig. 15).
In addition, there are two outstanding components of B-fields toward the regions characterized by the highest H2 column densities: one component associated with the peak of continuum emission at 154 m runs along the northeast to southwest direction with the median B-field orientation of 70; the other is located at the north of the first component and associated with the peak of continuum emission at 89 m, which exhibits the B-field vectors along the east-west direction with a median position angle of 88° and 81° at 89 and 154 m, respectively (see Fig. 15). We will further discuss the B-field morphology together with B-field strengths in Sect. 4.1.
@@ 3.2 Measurement of magnetic field strength using the DCF method
We first used the classical DCF method (Davis, 1951; Chandrasekhar & Fermi, 1953) (hereafter we refer to as the DCF method) to calculate the strengths of the B-field projected in the plane-of-the-sky () in the entire region of the Oph-A cloud. We compared our obtained results with those obtained using the more advanced techniques described in Sect. 3.3.
@@
where is the correction factor, is the gas mass density in units of g cm-3, =2.8 is the mean molecular weight per hydrogen molecule (Kauffmann et al., 2008), and is the mass of the hydrogen atom, is the gas volume density in units of cm-3. Then, is the one-dimensional (1D) non-thermal velocity dispersion along the line of sight (LOS) of the ideal gas species that traces the dense region, similarly to the dust-polarized emission in units of km s-1, and is the polarization angle dispersion in degrees. The correction factor, , is uncertain, and numerical simulations by Ostriker et al. (2001) show when the polarization angle dispersion does not exceed 25. However, varying values of will result in different strengths (see e.g., Skalidis et al., 2021; Liu et al., 2021, for more detailed discussions).
Here, we use the formula given in Crutcher (2004):
@@ 3.2.1 Polarization angle dispersion maps
It is essential to see how the B-field orientation changes locally within the molecular cloud. To do so, we calculated across the Oph-A region. We used the method presented in Hwang et al. (2021), for which is calculated pixel-by-pixel throughout the map of Oph-A. We summarize the procedure of calculating toward the pixel in the map as follows:
@@
We used a box covering the pixel with a size of 99 pixels, corresponding to two beam sizes of the SOFIA/HAWC observations (i.e., 15.6 and 36.4 in bands C and D, respectively).
@@
We calculated the mean value of polarization angles within the small box, , which implies the local mean polarization angle toward the pixel.
@@
For each pixel within the small box, we calculated the difference between the observed polarization angle and the mean value. We computed the root mean square (RMS) of the distribution of angle differences in the box. This RMS value indicates the polarization angle dispersion calculated toward the pixel.
@@
By repeating the process for all pixels in the entire region, we obtained the map of in Oph-A. We note that if there are more than 50 of pixels undefined (i.e., having NaN value of ) in the computing box, the result of polarization angle dispersion calculation toward pixel is ignored. We applied this criterion to keep the high quality of the estimate. It is also necessary to de-bias the polarization angle dispersion due to the uncertainty of the observed polarized angles.
Figure 3 shows the maps of polarization angle dispersion toward Oph-A using HAWC data in two bands C (left panel) and D (right panel), respectively. The polarization angle dispersion calculated using data in bands C and D cover ranges of 2.3–32.2 and 1.3–31.5, which correspond to the median values of 5.7° and 5.8°, respectively. The differences of in two wavelengths imply that the data at the shorter wavelength tends to trace warm dust surrounding the high-mass star S1 and toward the outer layers; whereas the data at the longer wavelength probes the colder dust in the denser layers of the cloud. We note, however, that in maps of in both bands, there are a few pixels where (27 pixels in band C and 45 pixels in band D). Thus, we do not use the value in these pixels to estimate in Sect. 3.2.4.
@@ 3.2.2 Velocity dispersion map
We used the HCO+(43) molecular data observed by JCMT/HARP to estimate the total velocity dispersion, . This molecular line has a critical density of 106 cm-3 (Greve et al., 2009) and thus is considered a good tracer for probing the dense region in the Oph-A cloud. The HCO+(43) emission is well-represented by a single Gaussian profile. Thus, in each pixel, we fit the line with a single Gaussian. The central velocity () across Oph-A covers the range from 2.25 to 3.76 km s-1 with a median value of 3.12 km s-1, which is in agreement with the systemic velocity of Oph-A of 3.44 km s-1 (André et al., 2007). We used the line width of the Gaussian as the measure of total velocity dispersion toward each pixel.
To construct the map of the 1D non-thermal (turbulent) velocity dispersion (), we subtracted the thermal component () from the map of the total velocity dispersion, given by:
where =29 amu is the mass of the HCO+ molecule, is the Boltzmann constant, and is the kinetic temperature of the gas. Here, we assume that the gas is in LTE condition and well-coupled with the dust in the cloud. Thus, the gas kinetic temperature is equal to the excitation temperature in all levels () and equal to the dust temperature: . We convolved the map of to the lower resolution of the molecular line map (14) to obtain the same pixel size on both maps. We converted the velocity dispersion of the non-thermal component into the FWHM velocity dispersion, given as . The uncertainty of non-thermal FWHMs is propagated from the uncertainty of the line width obtained from the Gaussian fit.
Figure 4 shows the map of the 1D non-thermal FWHM of the HCO+ molecular line toward Oph-A. The values are in the range from 0.62 to 2.98 km s-1 with a median value of 1.19 km s-1. We find a good agreement between our result obtained with HCO+(43) and those from other line tracers, namely, 0.95 km s-1 using N2H+(32) from APEX and 1.5 km s-1 using C18O(32) from JCMT/HARP (see Appendix C for details). Thus, it is plausible to use results from the HCO+(43) line in the following analysis.
@@ 3.2.3 Gas volume density map
We estimated the gas volume, , across the dense cloud from . Here, we assume that the cloud has a spherical Plummer-like shape (Liseau et al., 2015; Santos et al., 2019), which is frequently used to describe the globular dense cores (Plummer, 1911). We consider a sphere centered at the peak position of , located at (RA, Dec)= (, ), which is also associated with the position of the starless core SM1 (the separation is much less than 1″). We took the radius of the cloud (0.074 pc, Santos et al., 2019). To describe the effect of stellar feedback from the high-mass Oph-S1 star, we further add a spherical shell with a radius of surrounding this star. We also assume that the positions of peak column density and the S1 star are located in the same plane perpendicular to the LOS. The top panels in Fig. 5 show the 3D schematic view of the cloud (left) and a cross-section on the -plane at a given position (right). The - and -axes perpendicular to each other are in the plane of the sky and the axis is along the LOS. The center of the sphere is located at . The depth of the cloud at each position () along axis is calculated by:
where and are the radius of horizontal circles perpendicular to the plane of sky at position (see top right panel in Fig. 5). We obtain the map of , shown in the bottom left panel in Fig. 5. The values of are in the range of . We then infer from using the formula:
The bottom right panel in Fig. 5 shows the spatial extent of across Oph-A, overlaid by the contours representing the continuum emission in band D (Stokes-). The map has an angular resolution of 11.4″and a pixel size of 3.2, similar to that of the map. The gas density spans from cm-3 and is highest at the center of the cloud. The pattern of distribution well follows the intensity map in band D, revealing the ridge shape of the dense cloud. Some pixels located at the circular boundary in the map show anomalously high densities; this is due to the very low depths caused by the spherical model, and thus we exclude these pixels from our strength measurement.
@@ 3.2.4 Magnetic field strength maps
We used the maps of , , and to measure in the entire region of Oph-A. All input maps were smoothed to the angular resolution of 14, corresponding to that of the map. We then reconstructed the maps of at 89 and 154 m toward Oph-A using Eq. 2, which are shown in Fig. 6.
The values vary from mG and from mG for data at 89 and 154 m, respectively. The highest strength is found at the center region of the cloud, associated with the emission peak at 154 m and also the position of the SM1 core. The median values are and mG333The uncertainties of these median values are solely of the estimated distribution across the cloud. , as measured using 89 and 154 m data, respectively. The uncertainty of is propagated based on the uncertainty of the velocity dispersion of the gas and polarization angle dispersion. Since map is inferred from resulted in SED fitting models (see Santos et al., 2019). In addition, the estimated depth of the cloud is based on the 3D geometrical shape of the cloud and also is uncertain. Thus, we did not take into account the uncertainty of the gas volume density in the total uncertainty of the B-field strengths. For this reason, the uncertainty of might be underestimated.
@@ 3.3 Measurement of magnetic field strength using other methods
Several new techniques have been proposed to improve the DCF method for the last decades. For instance, Hildebrand et al. (2009) introduced the angle dispersion function (ADF) method where the dispersion in polarization angles in Eq. 2 is replaced by the square root of the structure function of polarization angles. This provides an improved measurement of B-field strength but also gives us the ratio of the turbulent to large-scale B-field strengths. Alternatively, Skalidis & Tassis (2021) introduced a modification of the DCF method (hereafter ST), assuming that the mean (ordered) and fluctuation components of the total B-field energy, and , respectively, are not independent. More recently, the differential measure approach (DMA; Lazarian et al., 2022) is considered promising to enhance the accuracy of B-field strength measurements at scales of smaller than the turbulence injection scale and takes into account the anisotropic properties of MHD turbulence (given by factor ). For the sake of cross-checking, the B-field strength measurement in Oph-A using the DCF method, we re-calculated the mean B-field strengths using the ADF, ST, and DMA methods at the two subregions (a) and (e) associated with the densest parts of the cloud (these subregions are named as in Kwon et al. (2018), see also in Fig. 15).